# Yesterday’s News: Benchmarking Multi-Dimensional Out-of-Distribution Generalisation of Misinformation Detection Models

### v1

@@ B.6 Publisher-level Metadata
    colspec=ccccccc, columns = colsep=4pt, cell33 = bg = lightgray, cell44 = bg = lightgray, cell55 = bg = lightgray, cell66 = bg = lightgray, cell77 = bg = lightgray, cell88 = bg = lightgray
    Table 13: Overlap of publishers between years.
@@
    Appendix C Publisher Statistics
    Across all dataset iterations, the vast majority of articles are written by a small minority of prolific publishers. This is shown in Figure 5. The most prolific publishers have authored between $40$-$60$% of articles within a year. The next deciles manage only $20$%, then $10$-$15$%, etc., with the smallest publishers authoring only a few articles. This effect seems to get more drastic for the later dataset years.
    As discussed in Sec. 5 this can obscure poor performance when using article-level evaluation metrics. In such cases, the average performance score will tend to be dominated by the performance of the largest publishers, instead of the body of publishers as a whole. As shown in Figure 1 and Tables 16 and 17, performance on these less mainstream publishers tends to be substantially worse. As a result, the OoD performance metrics in Table 2 likely underestimate the generalisation gap.
    Another noteworthy statistic is the amount of overlap between publishers across the dataset years. While the dataset as a whole contains $488$ distinct publishers, as can be seen in Table 7, each dataset year contains fewer total publishers. Furthermore, the distribution of those publishers (in terms of article counts) shifts across the years. This is partially due to a changing collection methodology used by Horne et al.; Norregaard et al.; Gruppi et al.; Gruppi et al.; Gruppi et al..
@@ Appendix D Training Details
    Table 14: An overview of important hyperparameters and their values. These were tuned on the validation set of the ‘Uniform’ split dataset.
@@
    Some other important training hyperparameters are listed in Table 14. We set a high learning rate for the classifier and pooler, but a very low value for the embeddings layer to avoid catastrophic forgetting. We employed a polynomial decaying learning rate scheduler, which first linearly warms up, before decaying according to η_t=(tmax_steps)^powerη_t-1.. For each experiment, regardless of the size of the dataset, we allow for a maximum of $63.0e+$ 6 update steps, evaluating on the validation set every $5$%. We employ early stopping with patience of 2, through which training usually concludes well before reaching the maximum budget.
@@ D.1 Additional Training Details
    All experiments were conducted on a Linux SLURM-based academic cluster. Nodes consisted of an Intel Xeon Platinum 8360Y CPU with 18 cores, 2.4 GHz, a single NVIDIA A100 GPU accelerator (40 GiB of HMB2 memory) and 128 GiB of DDR4 memory.
@@
    Figure 6: Probabilities of making an error for publishers of a particular label and political bias, relative to the probability of making a mistake for publishers of the same label. FP columns denote the probability of error for reliable publishers, and FN vice versa. The higher
    Appendix E Additional Results
    E.1 LLM Inference
    To conduct LLM inference, we use meta-llama/Meta-Llama-3-8B-Instruct using the following prompt:
    <s>[INST] <<SYS>> You are a content moderator working with journalistic articles. Your task is to identify articles from unreliable publishers. <</SYS>>
    Does the following text come from a reliable news publisher? Respond with ’yes’ or ’no’.
@@ E.1 LLM Inference
    E.2 Political Bias
@@
    $\frac{1-\mathbb{E}[\text{{Acc}}_{p}|\text{bias}(p),y_{p}]}{1-\mathbb{E}[\text{{Acc}}_{p}|y_{p}]}-1,$
@@
    $\exp\quad$ $\text{pmi}\left(\hat{y}_{p}\not=y;\text{bias}(p)|y\right)-$
    =$\frac{p(\hat{y}_{p}\not=y|\text{bias}(p),y)}{p(\hat{y}_{p}\not=y|y)}-1.$
    The higher this value, the greater the association between making an error for a particular political bias deviates from the expected value. If an error is made for a reliable publisher, this corresponds to a False Positive (FP) whereas an error for an unreliable publisher would correspond to a False Negative (FN). We introduce the shifting to ensure no bias corresponds with 0, whereas positive and negative bias correspond to positive and negative values, respectively.
    In Figure 6 we display various such biases for each iteration of the corpus. As seen in Table 16, performance degrades for the more extreme political biases. There is, however, a notable difference between left-biased and right-biased publishers, and especially for the reliable right-biased publishers. The probability of the model generating a False Positive is substantially higher (sometimes near twice as likely) than for reliable leftist or centre biased publishers. We conclude that the model tends to confuse articles from reliable right biased publishers with unreliable ones, far beyond what would be considered ’due to chance’.
@@
    E.3 Publisher Heterogeneity
    In Sec. 6.2 we tested for the effect of publisher heterogeneity on generalising to unseen publishers. We operationalized this by re-running the ‘Publisher’ split experiment with only the top-n most prolific publishers for each MBFC label, while leaving the test set untouched.
    This dramatically reduces the amount of variation in publishers, while minimally affecting the total amount of data present. This can be seen in Table 15. Note that these are expressed as percentages, but in absolute terms constitute hundreds of thousands of articles.
    Quantity Baseline Top 1 Top 2 Top 3
    Articles 100% 28% 43% 55%
    Publisher 100% 11% 20% 30%
    Table 15: Mean number of articles and publishers retained in the training set after limiting to the top-n most prolific publishers. Note that the number of articles and publishers deviates from the distributions depicted in Fig. 5. In this case, we additionally condition on the publisher’s MBFC label, across which publishers are not equally productive (see Table 7).
    E.4 Determinants of Publisher-Level Accuracy
@@
    $\text{{Acc}}_{p}=\frac{1}{|\mathcal{D}^{(p)}|}\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)}),$
@@
    $\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)})+\mathds{1}(\hat{y}^{(d)}\not=y^{(p)})$
    $=\sigma(\beta_{0}+\sum_{i=1}^{K}\beta_{i}x_{i}),$
@@
    E.4.1 Uniform Generalisation
    The fully specified model is displayed in table 16, with some important variables’ coefficients depicted graphically in Fig. 1. The left most column provides natural groupings of the different variables. The interpretation of the coefficient magnitudes should be adjusted to consider the range of possible values for the corresponding variable.
@@ E.4.1 Uniform Generalisation
    E.4.2 Topic Distance
    We repeat this analysis, but this time aggregating entities at the event/label level, using the ‘Topic’ split results. The model specification is left as above, with the introduction of a single additional variable: the minimum cosine distance to any event in the training set. The coefficients are provided in Table 17.
    While comparison across Tables 16 and 17 is not directly possible, as each estimates using different entities and different models, it is encouraging to see similar magnitudes for most variables’ coefficients.
@@ E.4.2 Topic Distance
    Table 16: Coefficients of the publisher-level determinants model, using the uniformly split models. Range indicates the possible values each variable can take, with ..absent... . indicating all natural numbers between the extremes (inclusive). $\beta$ provides the logistic parameter, and $\exp(\beta)$ its exponent (i.e. the log-odds ratio). Column ’Std. Err.’ provides the standard error, ’p’ the corresponding $p$-value and ’95% CI’ the confidence interval. In the parameter names, the ’:’ indicates an interaction term.
@@
    Table 17: Idem, but now for the publisher-topic aggregated binomial logistic model. Since the aggregation level differs from that of the model presented in Table 16, the coefficients are not directly comparable.

### v2

@@ B.7 Article Statistics
colspec=clccccc, hline1, 3, 13 = 1pt,solid, hline2 = 5,60.5pt, solid, hline9 = 0.5pt, dashed, cell11 = r=2,c=1c, cell12 = r=2,c=1c, cell13 = r=2,c=2c, cell15 = r=1,c=2c, cell17 = r=2,c=1c, # & Description # Articles Labels ID/OoD
reliable+ unreliable-
7 US Federal Elections 192k 18.8% 64.11% 35.89% ID
9 International Affairs 161k 15.8% 75.80% 24.20% ID
8 Health & COVID 149k 14.6% 70.85% 29.16% ID
4 Entertainment & Sports 112k 11.0% 86.30% 13.70% ID
5 Economy & Social Issues 112k 10.9% 63.50% 36.50% ID
6 Science & Technology 89k 8.7% 71.68% 28.31% ID
3 US Local Politics 66k 6.5% 68.85% 31.15% OoD
2 Crime & Justice 64k 6.2% 68.93% 31.07% OoD
0 Conflict 45k 4.4% 69.37% 30.63% OoD
1 Biden Administration 28k 2.7% 60.43% 39.57% OoD
B.8 Publisher Statistics
Across all dataset iterations, the vast majority of articles are written by a small minority of prolific publishers. This is shown in Figure B.2. The most prolific publishers have authored between $40$-$60$% of articles within a year. The next deciles manage only $20$%, then $10$-$15$%, etc., with the smallest publishers authoring only a few articles. This effect seems to get more drastic for the later dataset years.
As discussed in Section 6 this can obscure poor performance when using article-level evaluation metrics. In such cases, the average performance score will tend to be dominated by the performance of the largest publishers, instead of the body of publishers as a whole. As shown in Figure 1 and Tables E.2 and E.3, performance on these less mainstream publishers tends to be substantially worse. As a result, the OoD performance metrics in Table 2 likely underestimate the generalisation gap.
Figure B.2: The authorship of articles in the dataset, aggregated into deciles. The right most column of each panel represents the top 10% most prolific publishers, the left most the bottom 90%.
Another noteworthy statistic is the amount of overlap between publishers across the dataset years. While the dataset as a whole contains $488$ distinct publishers, as can be seen in Table B.6, each dataset year contains fewer total publishers. Furthermore, the distribution of those publishers (in terms of article counts) shifts across the years. This is partially due to a changing collection methodology used by Horne et al.; Norregaard, Horne, and Adali; Gruppi, Horne, and Adalı; Gruppi, Horne, and Adalı; Gruppi, Horne, and Adalı.
{tblr}
@@ B.8 Publisher Statistics
2017 2018 2019 2020 2021 2022
[…]
2018 0.43 1.00 0.62 0.34 0.33 0.32
2019 0.31 0.62 1.00 0.52 0.50 0.49
2020 0.17 0.34 0.52 1.00 0.86 0.80
2021 0.16 0.33 0.50 0.86 1.00 0.89
2022 0.16 0.32 0.49 0.80 0.89 1.00
Table B.8: Overlap of publishers between years.
@@
Appendix C Additional Information on Generalization Axes and Splits
Uniform
For this generalization axis, we employ the industry standard method of stratified and shuffled train/validation/test splits. We maintain the ratios 80/10/20, respectively, throughout. This is meant to form the baseline set of values, both in terms of ID-OoD articles performance, but also the expected delta between those.
Time
Misinformation changes quickly, but models have been found to be brittle to articles outside the time-span of the training data Bozarth and Budak (2020); Horne, Nørregaard, and Adali (2020). We train models on a dataset from a single year, and evaluate performance on all other years. To avoid conflating ‘Time’ and ‘Publisher’ results, we only evaluate on articles from publishers seen in the training set. The proportion of publishers–and as a result, the reliable/unreliable proportions–does change drastically across dataset years.
Event
We define events as occurrences that span news articles with a definite time-span and sudden occurrence. Novel events spawn articles with a markedly different vocabulary, introducing words, names, or terms not yet available to the model. Prior work has shown models suffer when evaluated on unseen events Lee et al. (2021); Ding et al. (2022). The dataset contains a multitude of such events, but we focus specifically on the COVID-19 pandemic; an event that particularly notable for its rapid onset and the volume of misinformation it spawned Bradd (2024). Models are trained on a subset of articles not containing COVID-19 keywords, and evaluated on ones that do. The article discovery-process is outlined in more detail in Appendix C.1.
Topic
We define topics to be relatively static collections of events, wherein the style or publishers’ opinions change little across years or decades. We discover such topics automatically, bottom-up. Specifically, we apply a heavily modified variant of the BERTopicGrootendorst (2022) algorithm to discover the many thousands of events that occur in each year of the dataset. Each event is textually represented by a TF-IDF vector. To group events into topics, we apply a second round of spectral clustering on the adjacency matrix induced by the inter-event cosine distance matrix. The number of articles contained by a topic differs substantially (see Table B.6). We reserve the $n$ smallest topics, which collectively contain roughly 20% of all articles for testing, and train on articles from all other topics.
Publisher
All scraped articles are annotated with the news website, outlet, or publisher that produced the article (see Figure B.1). In general, news publishers have some editorial bias or stance, making their corpus of articles distinctly different from that of other publishers. Prior work has found misinformation classifiers to be particularly sensitive to changes in publisher Barrón-Cedeño et al. (2019); Bozarth and Budak (2020). We specifically test for this by reserving the $n$ smallest publishers, which collectively contain roughly 20% of all articles for the test set, and train on articles from all other publishers. We stratify this splitting across political bias, to ensure each political bias is represented equally in both the train and test sets.
Political Bias
Separate from editorial bias, publishers tend to exhibit a political bias as well. MBFC has annotated all publishers on a left-right political bias continuum. We map all ‘Extreme Left’ and ‘Left’ publishers to a coarser ‘Left’ group, and vice versa for the ‘Right’-side of the political spectrum. All other political biases (‘Center-Left’, ‘Center-Right’, ‘Least Biased’ and ‘Pro-Science’) were mapped to ‘Center’ bias. We reserve either the ‘Left’ or ‘Right’ group for evaluation, and train on the collection of articles from the opposite political bias, along with all ‘Center’ biased publishers.
Misinformation Type
Misinformation presents in various forms. MBFC divides unreliable publishers in three categories: ‘Questionable-Source’ (publishers that exhibit extreme bias or propaganda), ‘Conspiracy-Pseudoscience’ (publisher aligning with known conspiracies or pseudo-scientific practices) and ‘Satire’ (publishers whose content is purposefully false for comedic effect). To test this form of generalization, we reserve either ‘Questionable-Source’ or ‘Conspiracy-Pseudoscience’ labels for evaluation, and train on all other articles. The test set includes a stratified sample of reliable sources as well.
C.1 Identifying COVID-19 Articles
Figure C.1: The top figure shows the proportion of articles published that day which discuss the COVID-19 pandemic. The solid line provides a LOESS smoothed trend line. The bottom figure provides a Google search interest for topics that fall under Coronavirus disease 2019. The different years (2020, 2021, 2022) are displayed as banded vertical columns.
For ‘Event’ splitting, we focus on a single event present over the latter 3 dataset years: the COVID-19 pandemic. We first derive a set of COVID-19 keywords (e.g., ‘sars-cov-2’, ‘lockdown’, ‘mask’), and combine it with the set of keywords defined by Gruppi, Horne, and Adalı (2021). We include all articles from 2020-2022 that include any of these terms in the held-out test set. While this is bound to induce a large number of false positives, it ensures no COVID-19 related terms contaminate the models’ learned vocabulary. In Figure C.1 we display the correlation between the found number of articles and the amount of Google search volume.
C.2 Test Set Correlations
Year IoU Random IoU Execess IoU Test Set Sizes
Publisher Topic
2017 0.1183 0.1084 9.07% 19764 20709
2018 0.1227 0.1162 5.62% 109043 85832
2019 0.1210 0.1133 6.80% 133465 110168
2020 0.1386 0.1116 24.24% 203224 205873
2021 0.1317 0.1055 24.82% 192315 203311
2022 0.1318 0.0992 32.93% 172530 186986
Table C.1: The amount of article_id overlap between the ‘Publisher’ and ‘Topic’ test sets. Column ‘IoU’ gives the intersection over the union between the two held-out test sets. Column ‘Over Random’ compares this to the achieved value for $10000$ simulated draws from a uniform distribution of the same sizes as the actual test set sizes (given in the last two columns).
Publishers have a tendency to prioritize different topics. This is especially prevalent for less mainstream publishers, which have far fewer resources to spend on the breadth of their reporting. It is therefore plausible that the ‘Publisher’ and ‘Topic’ splits have some overlap.
We test for this by comparing the amount of article level overlap in the held-out test sets of the ‘Publisher’ and ‘Topic’ to random draws of the same size from uncorrelated uniform distributions. The larger this overlap, the more correlated the two test sets are. We display the results of our test in Table C.1. While we find more overlap than random draws, the effect is not large. Even in the most egregious case, there is only a $0.04$ difference in absolute IoU. As such, we believe the two different test sets measure distinct generalization aspects.
@@ Appendix D Training Details
Table D.1: An overview of important hyperparameters and their values. These were tuned on the validation set of the ‘Uniform’ split dataset.
@@
Some other important training hyperparameters are listed in Table D.1. We set a high learning rate for the classifier and pooler, but a very low value for the embeddings layer to avoid catastrophic forgetting. We employed a polynomial decaying learning rate scheduler, which first linearly warms up, before decaying according to
$\eta_{t}=\left(\frac{t}{\text{{max\_steps}}}\right)^{\texttt{power}}\eta_{t-1}$
For each experiment, regardless of the size of the dataset, we allow for a maximum of $63.0e+$ 6 update steps, evaluating on the validation set every $5$%. We employ early stopping with patience of 2, through which training usually concludes well before reaching the maximum budget.
@@ D.1 Additional Training Details
All experiments were conducted on SNELLIUS, a Linux SLURM-based supercomputer. Nodes consist of an Intel Xeon Platinum 8360Y CPU with 18 cores, 2.4 GHz speed, and a single NVIDIA A100 GPU accelerator (40 GiB of HMB2 memory), and 128 GiB of DDR4 memory.
@@
D.2 LLM Inference
To conduct LLM inference, we use Meta’s llama-3-8B-Instruct using the following prompt:
<<SYS>>
You are a content moderator working with journalistic articles. Your task is to identify articles from unreliable publishers.
<</SYS>>
Does the following text come from a reliable news publisher?
Respond with ’yes’ or ’no’.
@@ D.2 LLM Inference
D.3 Annotating Article Properties
Subjectivity
{tblr}
colspec=ccccccc, columns = colsep=1pt, vline3 = 3-80.5pt, solid, hline3 = 3-80.5pt, solid, cell33 = bg = lightgray, cell44 = bg = lightgray, cell55 = bg = lightgray, cell66 = bg = lightgray, cell77 = bg = lightgray, & E. Obj. M. Obj. Mixed M. Subj. E. Subj.
E. Obj. 5.71% 5.13% 0.39% 0.01%
M. Obj. 1.80% 12.30% 3.76% 0.97%
Mixed 3.37% 15.19% 7.62% 0.49%
M. Subj. 0.05% 0.39% 4.35% 14.70% 4.00%
E. Subj. 0.83% 5.08% 13.77%
Table D.2: A confusion matrix between the subjectivity assessments of ChatGPT4o (rows) and ChatGPT4o-mini (columns).
To annotate articles for subjectivity, we used ChatGPT4o-mini, and ChatGPT4o. The former was used to annotate the bulk of articles, and the latter to verify the overall quality of the annotations. We used the following prompt:
<<SYS>>
You are a helpful assistant, helping analyse the properties of news articles. Before a final answer, make sure to explain your thinking.
<</SYS>>
Please classify how objective the following article is.
Objective articles take a neutral stance on topics, and focus on reporting factual news.
Subjective articles instead focus on opinions, which are more difficult to verify and can take specific stances for or against topics.
The title and body are provided. After you provide your reasoning, respond with one of entirely objective, mostly objective, mixed, mostly subjective, entirely subjective, and nothing else.
Article: \${ARTICLE}
Table D.2 gives the classification correspondence between the two models. Overall, we find that the mini model aligns well with the larger model, and that the larger model tends towards more subjective annotations, rather than less. As such, it is plausible that the subjectivity within the dataset remains underreported. We include some examples of articles and their subjectivity annotations in https://github.com/ioverho/misinfo-general/tree/main/assets/subjectivity_annotations_examples.
Emotion
We perform a similar analysis, but now for emotions present in the articles. We use the following prompt, and only use ChatGPT4o-mini:
<<SYS>>
You are a helpful assistant, analysing the properties of news articles. Before a final answer, make sure to explain your thinking.
<</SYS>>
Please identify the dominant emotions in the following article.
We will be using Plutchik’s 8 basic emotions, along with a label for neutral:
  1. 1.
    Joy: a feeling of happiness, pleasure, or contentment.
  2. 2.
    Sadness: a feeling of loss, disappointment, or grief
  3. 3.
    Trust: a sense of safety, security, and connection with others
  4. 4.
    Disgust: a strong aversion to something unpleasant, often related to taste, smell, or moral judgments
  5. 5.
    Fear: a response to perceived danger, leading to caution or escape behaviors
  6. 6.
    Anger: a reaction to perceived threats or injustice
  7. 7.
    Surprise: a reaction to something unexpected
  8. 8.
    Anticipation: looking forward to or expecting something, which can bring excitement or anxiety
  9. 9.
    Neutral: emotional balance, no strong positive or negative emotions are present
To respond, please list the emotions present in the article. When labelling the article for the ’Neutral’ emotion, please make sure no other emotion is present.
The article’s title and body are provided. After you provide your reasoning, respond with a list of joy, sadness, trust, disgust, fear, anger, surprise, anticipation, neutral, and nothing else.
Article: \${ARTICLE}
Figure D.1: Emotion propensity, like Figure 5, but now in absolute values, and presented for each emotion in isolation. More intense, red colors signify a higher presence of an emotion, whereas cooler, blue colors the opposite.
Figure D.1 presents the same data as in Figure 5, but now transposed (all publisher class combinations per emotion), and in absolute values. We find emotion overall to be present in all forms of articles and publishers.
Appendix E Additional Results
E.1 Political Bias
Figure E.1: Probabilities of making an error for publishers of a particular label and political bias, relative to the probability of making a mistake for publishers of the same label. FP columns denote the probability of error for reliable publishers, and FN vice versa. The higher
@@ E.1 Political Bias
$\frac{1-\mathbb{E}[\text{{Acc}}_{p}|\text{bias}(p),y_{p}]}{1-\mathbb{E}[\text{{Acc}}_{p}|y_{p}]}-1$
@@
$\frac{p(\hat{y}_{p}\not=y|\text{bias}(p),y)}{p(\hat{y}_{p}\not=y|y)}-1=\exp\left(\text{pmi}\left(\hat{y}_{p}\not=y;\text{bias}(p)|y\right)\right)-1$
The higher this value is, the greater the association between making an error for a particular political bias deviates from the expected value. If an error is made for a reliable publisher, this corresponds to a False Positive (FP) whereas an error for an unreliable publisher would correspond to a False Negative (FN). We introduce the shifting to ensure no bias corresponds with 0, whereas positive and negative bias correspond to positive and negative values, respectively.
In Figure E.1 we display various such biases for each iteration of the corpus. As seen in Table E.2, performance degrades for the more extreme political biases. There is, however, a notable difference between left-biased and right-biased publishers, and especially for the reliable right-biased publishers. The probability of the model generating a False Positive is substantially higher (sometimes near twice as likely) than for reliable leftist or centre biased publishers. We conclude that the model tends to confuse articles from reliable right biased publishers with unreliable ones, far beyond what would be considered ’due to chance’.
@@
E.2 Publisher Heterogeneity
Quantity Top 1 Top 2 Top 3
Articles 28% 43% 55%
Publisher 11% 20% 30%
Table E.1: Mean number of articles and publishers retained in the training set after limiting to the top-n most prolific publishers. Note that the number of articles and publishers deviates from the distributions depicted in Figure B.2. In this case, we additionally condition on the publisher’s MBFC label.
In Section 7.2 we tested for the effect of publisher heterogeneity on generalising to unseen publishers. We operationalized this by re-running the ‘Publisher’ split experiment with only the top-n most prolific publishers for each MBFC label, while leaving the test set untouched.
This dramatically reduces the amount of variation in publishers, while minimally affecting the total amount of data present. This can be seen in Table E.1. Note that these are expressed as percentages, but in absolute terms constitute hundreds of thousands of articles.
E.3 Determinants of Publisher-Level Accuracy
@@
$\text{{Acc}}_{p}=\frac{1}{|\mathcal{D}^{(p)}|}\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)})$
@@
$\sum_{d\in\mathcal{D}^{(p)}}\mathds{1}(\hat{y}^{(d)}=y^{(p)})+\mathds{1}(\hat{y}^{(d)}\not=y^{(p)})\\ =\sigma(\beta_{0}+\sum_{i=1}^{K}\beta_{i}x_{i})$
@@
Uniform Generalisation
The fully specified model is displayed in table E.2, with some important variables’ coefficients depicted graphically in Figure 1. The left most column provides natural groupings of the different variables. The interpretation of the coefficient magnitudes should be adjusted to consider the range of possible values for the corresponding variable.
@@
Topic Distance
We repeat this analysis, but this time aggregating entities at the event/label level, using the ‘Topic’ split results. The model specification is left as above, with the introduction of a single additional variable: the minimum cosine distance to any event in the training set. The coefficients are provided in Table E.3.
While comparison across Tables E.2 and E.3 is not directly possible, as each estimates using different entities and different models, it is encouraging to see similar magnitudes for most variables’ coefficients.
@@
Table E.2: Coefficients of the publisher-level determinants model, using the uniformly split models. Range indicates the possible values each variable can take, with ..absent... . indicating all natural numbers between the extremes (inclusive). $\beta$ provides the logistic parameter, and $\exp(\beta)$ its exponent (i.e. the log-odds ratio). Column ’Std. Err.’ provides the standard error, ’p’ the corresponding $p$-value and ’95% CI’ the confidence interval. In the parameter names, the ’:’ indicates an interaction term.
@@
Table E.3: Idem, but now for the publisher-topic aggregated binomial logistic model. Since the aggregation level differs from that of the model presented in Table E.2, the coefficients are not directly comparable.
Acknowledgements.
[…]